OpenStudy (anonymous):

Which number correctly completes this equation? 6 --- % * 45 = _____ 10 Select one of the options below as your answer: A. .027 B. .27 C. 2.7 D. 27

OpenStudy (anonymous):

\[\Huge \frac{6}{10}\%\times 45=...\] ?

OpenStudy (anonymous):

any percent number like x% in fraction will be \(\frac{x}{100}\) so I guess you have there... \[\Huge \frac{\frac{6}{10}}{100}\cdot 45\] ...

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

= 6 * 45 /1000. = 27/100

OpenStudy (anonymous):

\[\Large \frac{6 }{10\cdot 100 }\cdot 45=\frac{6}{10000}\cdot 45=\] \[\LARGE =0.027\]

OpenStudy (anonymous):

ohh sorry... I added a zero there... let me fix it :$

OpenStudy (anonymous):

I Think it's C

OpenStudy (anonymous):

it's \[\LARGE 0.27=\frac{6*45}{1000}\]

OpenStudy (anonymous):

thank you.